molarity of aqueous solution containing 129 g of glucose(C6H12O6) in 200 mL of solution is
1
Expert's answer
2020-04-20T15:06:29-0400

"molarity=\text{moles per litre}"

Number of moles=129/180= 0.717moles

molarity=0.717/200ml= 0.0036M